How many quarts of pure antifreeze must be added to 7 quarts of a 50% antifreeze solution to obtain a 60% antifreeze solution?
Thanks!

Answer by stanbon(75887)

How many quarts of pure antifreeze must be added to 7 quarts of a 50% antifreeze solution to obtain a 60% antifreeze solution?
------
Equation:
alcohol + alcohol = alcohol
0.50*7 + x = 0.60*(7+x)
---
Multiply thru by 100 to get:
50*7 + 100x = 60*7 + 60x
----
40x = 10*7
x = (1/4)7
x = 1.75 quarts (amt. of pure antifreeze needed)
=================
Cheers,
Stan H.
===================